A man using a 70kg garden roller on a level surface exerts a force of 200 N at 45 degree to the ground.What is the vertical force of the roller on the ground,if he pushes the roller?
  • 70N
  • 560N
  • 840N
  • 200N

Grade:12th Pass

1 Answers

Arun
6 years ago
Dear Ashwin
 
The force of the roller is the weight plus the vertical component of the gardener force 

W = 70kg * 10m/s^2 = 700 N 
G = +/- 200N * sin(45) = +/- 141 N 

When he pulls, G is negative and 

F = 700N - 141N = 559 N 

When he push, G is positive and 

F = 700N +141N = 841 N
 
 
Regards
Arun (askIITians forum expert)